OpenStudy (anonymous):

verify that the following functions are inverses of each other. 1. 3x+4y=6 y=-4/3 x +2 i need help..

OpenStudy (aum):

3x + 4y = 6 4y = -3x + 6 y = -3/4 * x + 6/4 y = -3/4 * x + 3/2 \(f(x) = -3/4x + 3/2\) If \(f^{-1}(x) = -4/3x + 2\), then \(f(f^{-1}(x)) = x\) Find \(f(f^{-1}(x)) \)